# Concurrent Balanced Augmented Trees

### v1

@@ 7. Experimental Results
We implemented BAT using the chromatic tree implementation in (Brown et al., 2014), which uses LLX/SCX primitives from (Brown et al., 2013), in the publicly available SetBench (UW Multicore Lab, 2025) microbenchmark. Our experiments show that (1) delegation can improve BAT’s insert and delete throughput by over 100%, (2) BAT scales well with thread count and data structure size, and (3) BAT performs significantly faster than previous data structures in workloads with more than 2% order-statistic queries and in workloads with range queries of size larger than 2K-10K.
Experiment Figure Description Purpose
Improvement Study 5, 5 Throughput vs number of threads on an update only workload Compare our variants against an unbalanced augmented tree under uniform and skewed workloads
Query Scalability 5 Throughput vs number of threads for queries on BAT-EagerDel Compare the scalability of different queries on our balanced augmented tree
Range Query Size 8 Throughput vs range query size for a 20% update workload Show the benefits of augmenting a concurrent tree
Size Scalability 10 Throughput vs data structure size for uniform and Zipfian distributions Show how various trees scale to larger data structure sizes
Table 1. Description of our experiments
(a) MK 10M, 50-50-0-0. Comparing our three variants.
(c) RQ 50K, MK 10M 5-5-0-90. Scalability of queries on BAT-EagerDel.
Figure 5. Performance of variants of our BAT.
Data Structures. The following table summarizes the key properties of the data structures we compare.
@@
All are linearizable, written in C++, integrated into SetBench, and use the same memory reclamation scheme. Data structures that do not maintain augmented values achieve linearizable range queries by taking a snapshot and iterating over the range. Each implementation (other than FR-BST, which we implemented) was taken from its original paper.
Among the three existing concurrent augmented trees (Kokorin et al., 2024; Sela and Petrank, 2024; Fatourou and Ruppert, 2024), we chose to compare with FR-BST (Fatourou and Ruppert, 2024) because it has the best theoretical bounds.  (Kokorin et al., 2024) has a Kotlin implementation that is several orders of magnitude slower than our C++ implementation of FR-BST, according to the results in their paper. This cannot fully be attributed to language differences. We are not aware of any implementations of (Sela and Petrank, 2024).
@@
Key Distribution: Either uniform, sorted or Zipfian with parameter 0.95. The sorted and Zipfian workloads result in high contention as updates are routed to the same parts of the tree. The distribution was uniform unless otherwise specified.
@@
(a) $T$ 120, RQ 50K, 25-25-25-25. Size scalability, high update Zipfian workload.
Figure 10. Comparison of BAT variants to other trees with respect to data structure size.
@@
Range Query Size. Figure 8 shows the performance of various structures under varying range query size. Figures 8 and 8 show results for small and large trees. Since the unaugmented trees perform work proportional to the number of keys in the range, their performance drops off sharply for larger range queries. In contrast, in the augmented trees (FR-BST and ours), queries only perform work proportional to the height of the BST, so their performance stays consistent no matter the range query size. VerlibBtree outperforms the other non-augmented trees since it uses higher fanout trees for better cache efficiency, but loses out to the augmented structures after range query sizes reach 2000–4000. After reaching range query size 2M for a tree of size 10M, BAT-EagerDel is  400x as fast as the closest non-augmented tree. However, the added overhead for inserts and deletes causes the augmented structures to lose out heavily when range queries only traverse a few keys. For range queries of only 8 keys, BAT-EagerDel is 15x slower than VerlibBtree. BAT-EagerDel is around 3x faster than FR-BST because balancing reduces the average depth of the leaves and delegation reduces contention at higher levels of the tree.
@@
Thread Scalability. Figures 8 and 8 show scalability under low (5%) and high (50%) update percentages, respectively. These update percentages were selected according to YCSB workloads A and B (Cooper et al., 2010). FR-BST scales less well since it has higher contention close to the root, making updates perform worse when more threads are involved. The scaling of the other data structures are similar to each other, however BAT-EagerDel outperforms the closest unaugmented competitor by around 4x on the high update workload and 30x on the low update workload for all thread counts.
Isolated Performance. Figure 10 shows the average latency in seconds for updates and range queries under the same workload as Figure  8. In the update graph, we see that the performance of inserts and deletes on BATremains relatively constant. Furthermore, BAT-EagerDel has a lower range query latency than all unaugmented competitors when at least 2000 elements are being queried.
Size Scalability and Zipfian Distribution. Figure 10 shows the effect of increasing the data structure size (by varying the maximum key). We include the BAT variant with no delegation in these graphs to show it is still worse than the versions with delegation in the case of a Zipfian distribution. Overall, we see that BAT-EagerDel scales slightly better with size compared to VerlibBtree, BundledCitrusTree and VcasBST. VerlibBtree performs  15% worse on the Zipfian distribution (Figure 10) for small trees, while the others stay relatively constant.
Why Balancing Improves Throughput. BAT performs extra work to balance the tree, and calls to propagate occasionally have to traverse backwards or fill in nil versions. Nevertheless, our results show that BAT variants consistently outperform FR-BST, even on uniform and Zipfian key distributions, where FR-BST would be fairly balanced. We provide some key statistics to explain this. We measured on a workload with 120 threads, 100K max key, 50K range query size and an even percentage of inserts, deletes, range queries and finds on both a uniform and Zipfian distribution with parameter 0.99. On the BAT variants, each propagate only traverses 6.4% (5.9%) more nodes beyond the initial search path for the uniform (Zipfian) distribution. A call to propagate fills in only 0.075 (0.03) nil versions on average. Lastly, the average number of CASes attempted during a propagate call is 22.2 (22.4) for BAT, 13.9 (13.2) for BAT-EagerDel and 26.8 (27.5) for FR-BST. Thus, the extra costs incurred by rebalancing are minimal compared to the advantages of maintaining a more carefully balanced tree.
@@ 8. Conclusion
Augmentation makes search trees significantly more versatile by extending the set interface to enable support for aggregation queries, order-statistic queries, and range queries. In this paper, we designed, implemented, and empirically validated BAT, the first lock-free Balanced Augmented Tree. While we emphasized our augmentation scheme as applied to a chromatic tree, our scheme is general—adaptable to concurrent search trees where updates modify the tree by replacing one patch by another patch of new nodes.
@@ Appendix B Correctness
We follow the arguments similar to those for augmented binary search trees in (Fatourou and Ruppert, 2024) and extend them for the augmented chromatic trees presented in this paper. We first present the proof for the non-delegating version.

### v2

@@ 7. Experimental Results
Table 1. Summary of data structure properties
@@
We implemented BAT using the chromatic tree implementation in (Brown et al., 2014), which uses LLX/SCX primitives from (Brown et al., 2013), in the publicly available SetBench (UW Multicore Lab, 2025) microbenchmark. Our experiments show that (1) delegation can improve BAT’s insert and delete throughput by up to 200%, (2) BAT scales well with thread count and data structure size, and (3) BAT performs significantly faster than previous data structures in workloads with more than 0.15-11% order-statistic queries (depending on data structure size) and in workloads with range queries of size larger than 2K-10K. Our code is available at https://github.com/evanwrench/CBAT.
Table 2. Description of our experiments
Experiment Figure Description Purpose
Improvement Study 5, 5 Throughput vs number of threads on an update only workload Compare variants of BAT with FR-BST under uniform and skewed workloads
Query Scalability 5 Throughput vs number of threads for rank, select and range queries on BAT-EagerDel Compare the scalability of different queries on our balanced augmented tree
Range Query Size 8 Throughput vs range query size for a 20% update workload Show the benefits and tradeoffs of augmenting a concurrent tree
Size Scalability 10 Throughput vs data structure size with keys chosen using a Zipfian distribution Show how various trees scale to larger data structure sizes
(a) MK 10M, 50-50-0-0. Comparing our three variants.
(c) RQ 50K, MK 10M 5-5-0-90. Scalability of queries on BAT-EagerDel.
Figure 5. Performance of variants of BAT.
Data Structures. The table 1 summarizes the key properties of the data structures we compare. All are linearizable, written in C++, integrated into SetBench, and use the same memory reclamation scheme. Data structures that do not maintain augmented values achieve linearizable range queries by taking a snapshot and iterating over the range. Each implementation (other than FR-BST, which we implemented) was taken from its original paper.
@@
Key Distribution: Either uniform, sorted or Zipfian. The sorted and Zipfian workloads result in high contention as updates are routed to the same parts of the tree. The distribution was uniform unless otherwise specified.
@@
Figure 10. Comparison of BAT variants to other trees with respect to data structure size. $T$ 120, RQ 50K, 25-25-25-25. Size scalability, high update Zipfian (parameter=0.95) workload.
@@
Range Query Size. Figure 8 shows the performance of various structures under varying range query size. Figures 8 and 8 show results for small and large trees. Since the unaugmented trees perform work proportional to the number of keys in the range, their performance drops off sharply for larger range queries. In contrast, in the augmented trees (FR-BST and ours), queries only perform work proportional to the height of the BST, so their performance stays consistent no matter the range query size. VerlibBtree outperforms the other non-augmented trees since it uses higher fanout trees for better cache efficiency, but loses out to BAT-EagerDel after range query sizes reach 2000–4000. After reaching range query size 2M for a tree of size 10M, BAT-EagerDel is  400x as fast as the closest non-augmented tree. However, the added overhead for inserts and deletes causes the augmented structures to lose out heavily when range queries only traverse a few keys. For range queries of only 8 keys, BAT-EagerDel is 15x slower than VerlibBtree. BAT-EagerDel is around 3x faster than FR-BST because balancing reduces the average depth of the leaves and delegation reduces contention at higher levels of the tree.
@@
Thread Scalability. Figures 8 and 8 show scalability under low (5%) and high (50%) update percentages, respectively. These update percentages were selected according to YCSB workloads A and B (Cooper et al., 2010). FR-BST scales less well than other data structures since it has higher contention close to the root, making updates perform worse when more threads are involved. The scaling of the other structures are similar to each other, however BAT-EagerDel outperforms the closest unaugmented competitor by around 4x on the high update workload and 30x on the low update workload for all threads.
Isolated Performance. Figure 10 shows the average latency in seconds for updates and range queries under the same workload as Figure  8. In the update graph, we see that the performance of inserts and deletes on BAT-EagerDel remains relatively constant. Furthermore, BAT-EagerDel has a lower range query latency than all unaugmented competitors when at least 2000 elements are included in the query.
Size Scalability and Zipfian Distribution. Figure 10 shows the effect of increasing the data structure size (by varying the maximum key). We include the BAT variant with no delegation in this graph to show that delegation still has benefits when keys are chosen from a Zipfian distribution. Overall, we see that BAT-EagerDel scales slightly better with size compared to VerlibBtree, BundledCitrusTree and VcasBST. We only show results for Zipfian distribution, as the results for the uniform distribution are nearly identical.
Why Balancing Improves Throughput. BAT performs extra work to balance the tree, and calls to propagate occasionally have to retraverse parts of the tree or fill in nil versions. Nevertheless, our results show that BAT variants consistently outperform FR-BST, even in workloads where FR-BST is balanced in expectation. We provide some key statistics to explain this. We measured on a workload with 120 threads, 100K max key, 50K range query size and an even percentage of inserts, deletes, range queries and finds on both a uniform and Zipfian distribution with parameter 0.99. On the BAT variants, each propagate only traverses around 6.4% (5.9%) more nodes beyond the initial search path for the uniform (Zipfian) distribution. A call to propagate fills in only 0.075 (0.03) nil versions on average. Lastly, the average number of CASes attempted during a propagate call is 22.2 (22.4) for BAT, 13.9 (13.2) for BAT-EagerDel and 26.8 (27.5) for FR-BST. Thus, the extra costs incurred by rebalancing are minimal compared to the advantages of maintaining a more carefully balanced tree.
@@ 8. Conclusion
Augmentation makes search trees significantly more versatile by extending the set interface to enable support for aggregation queries, order-statistic queries, and range queries. In this paper, we designed, implemented, and empirically validated BAT, the first lock-free Balanced Augmented Tree supporting generic augmentation functions. While we emphasized our augmentation scheme as applied to a chromatic tree, our scheme is general—adaptable to concurrent search trees where updates modify the tree by replacing one patch by another patch of new nodes.
@@ Appendix B Correctness
We follow the arguments similar to those for augmented binary search trees in (Fatourou and Ruppert, 2024) and extend them for the augmented chromatic trees presented in this paper. We first present the proof for the non delegating version.